### v2

# A projection method for particle resampling

Abstract
Particle discretizations of partial differential equations are advantageous for high-dimensional kinetic models in phase space due to their better scalability than continuum approaches with respect to dimension. Complex processes collectively referred to as particle noise hamper long-time simulations with particle methods. One approach to address this problem is particle mesh adaptivity or remapping, known as particle resampling. This paper introduces a resampling method that projects particles to and from a (finite element) function space. The method is simple; using standard sparse linear algebra and finite element techniques, it can adapt to almost any set of new particle locations and preserves all moments up to the order of polynomial represented exactly by the continuum function space.
This work is motivated by the Vlasov-Maxwell-Landau model of magnetized plasmas with up to six dimensions, $3X$ in physical space and $3V$ in velocity space, and is developed in the context of a $1X$ + $1V$ Vlasov-Poisson model of Landau damping with logically regular particle and continuum phase space grids. The evaluation codes are publicly available, along with the data and reproducibility artifacts, and developed in the PETSc numerical library (petsc.org).
1 Background
Particle, marker particle, or macro-particle methods, such as particle-in-cell (PIC), are discretizations that, akin to traditional continuum-based methods such as finite elements (FE), finite volume, etc., discretize continuous PDE models as opposed to discrete, ground truth models like molecular dynamics. Particle methods scale with dimension with a theoretical accuracy of $\mathcal{O}(N^{-\frac{1}{2}})$ and complexity $\mathcal{O}(N)$, with $N$ particles, whereas continuum methods have higher order accuracy, $\mathcal{O}(N^{-p})$ for some order $p$, usually two or higher, the complexity is order $\mathcal{O}(N^{D})$ with $N$ grid points in each dimension $D$. For example, with a commonly attainable $p=2$ the complexities cross over at $D=4$ and particle methods have lower order complexity at higher dimensions. This scaling with dimension has motivated the use of the PIC methods for the Vlasov-Maxwell-Landau (VML) system, or Boltzmann’s equation with Coulomb collisions in Landau form.
Mesh or grid adaptivity is a fundamental tool in PDE modeling for both continuum and particle grid methods. Particle adaptivity is known as particle resampling in the physics community has been developed by many groups. Lapenta developed a method in the 1990s with a solve of the form $\left(M_{p}M_{p}^{T}\right)^{-1}$ for Lagrange multipliers that enforce moment constraints explicitly [24, 23, 25], that is formally similar to the pseudoinverse solve in the projection of our approach. Colella et al., developed a particle remapping method in a finite volume context that is similar to our approach with a direct remap from the grid to particles [41, 30, 31]. Faghihi et al. developed resampling methods with moment constraints and linear programming to enforce moments and other algebraic constraints [10]. Gonoskov proposed probabilistic down-sampling algorithms using algebraic constraints to enforced conservation [13]. Pfeiffer et al. introduced two conservative particle split and merge methods that use statistical properties of the plasma such as thermal speed [32]. Several groups have presented particle coalescence and splitting schemes, often using trees, on small groups of locally binned particles [37, 40, 4, 43, 26].
This paper develops a particle resampling approach that starts with a conservative mapping between particles and continuum grids, a projection [34], that conserves an arbitrary number of moments exactly, inspired by the structure preserving discretization community [20, 42]. The focus of this paper is to investigate the properties of this projection technique with a static regular remapping grids for both particles and the finite element space and a standard plasma model problem. This method allows for remapping a particle distribution to any new set of particles, supporting adaptivity in both the continuum grid construction and the particle grid definition, which is the subject of future work. The testing codes are built on PETSc (Portable, Extensible Toolkit for Scientific Computation), and are publicly available (Appendix §A).
This paper proceeds with relevant background in structure preserving methods in §2, the projection based resampling method in §3, numerical methods and test problem in §4, §5 experiments with the direct remap method, with a finite element version, experiments with the full high-order finite element projection and remapping method are presented in §6, and §7 concludes with a discussion of future work with this projection resampling method.
2 Structure preserving methods for Boltzmann’s equations
The critical new idea in this work comes from research on structure preserving methods for Boltzmann’s equations in general and the VML system for magnetized plasmas in particular, that results in a simple and elegant algorithm that provably conserve arbitrary number of moments.
Hamiltonian models in phase space where density is a function of both space ($\mathbf{x}$) and velocity or momentum space ($\mathbf{v}$) are the fundamental equation of gravitational dynamics and electrostatics plasmas, the Vlasov-Poisson system, and electromagnetic plasmas, the Vlasov-Maxwell system. A coulomb collision term accounts for the statistics of particle interactions not present in the Hamiltonian [14, 16, 3, 39, 18], giving rise to the governing equations for magnetized plasmas where the density of each species $\alpha$ is evolved in phase space according to
$\frac{df_{\alpha}}{dt}\equiv\frac{\partial f_{\alpha}}{\partial t}+\frac{\partial\vec{x}}{\partial t}\cdot\nabla_{x}f_{\alpha}+\frac{\partial\vec{v}}{\partial t}\cdot\nabla_{v}f_{\alpha}=\sum_{\beta}C\left[f_{\alpha},f_{\beta}\right]_{\alpha\beta},$
where the collisional term is summed over all species $\beta$. This equation is composed of the symplectic Vlasov-Maxwell term $\frac{df}{dt}=0$ and a metric, or diffusive, collision operator $C$, and Maxwells’s equations provide an expression for $\frac{\partial\vec{v}}{\partial t}=a=\frac{q_{\alpha}}{m_{\alpha}}\left(\mathbf{E}+\mathbf{v}\times\mathbf{B}\right)$.
This system has rich mathematical structure that can be preserved with proper discretizations. The metriplectic formalism is an approach to analyze VML and to develop structure preserving discretizations for the VML sysetm [16, 20, 15]. When a structure preserving grid based collision operator [16, 2, 1, 3] is coupled with a PIC method, a mechanism is needed to map distribution functions, in velocity space, between a particle representation and a finite element basis representation of the distribution function that preserves moments, as well as other structure [34]. Preserving the second moment in velocity space, energy, is critical for many applications. This mapping mechanism is central to the algorithms developed herein.
2.1 Structure preserving particle-finite element basis mapping
To apply a continuum operator in a PIC method that conserves moments a conservative particle-finite element basis mapping and remapping method is required. Given a particle with weight $w_{p}$ and position $\textbf{x}_{p}$ and a delta function representation $f_{p}(\textbf{x})=w_{p}\delta(\textbf{x}-\textbf{x}_{p})$. Given a finite element (FE) space $V$ of functions $\phi_{i}$ and coefficients $\rho_{i}$, a function can be expressed as $f_{FE}(\textbf{x})=\sum_{i}\rho_{i}\phi_{i}(\textbf{x})$. Ideally $f_{FE}(\textbf{x})=f_{p}(\textbf{x})$, but that not possible. Weak equivalence can however be enforced with:
[…]
$M\rho=M_{p}w,$
which defines an equation for particle deposition:
(2) $\rho\leftarrow M^{-1}M_{p}w.$
This mapping is proven to conserve moments up to the order polynomial that the FE space can represent exactly [17, 34], for example a quadratic element mesh is sufficient to conserve energy in velocity space.
After deposition on the FE space, a Poisson or Ampere’s law solve can be executed or the collision operator, $L$, can be evolved, $u\leftarrow L\rho$. In mapping $u$ back to particles one can simply invert (2), $\overline{w}\leftarrow M_{p}^{-1}Mu$, however $M_{p}$ is rectangular in general. The key idea from the structure preserving literature, where it is used to prove conservation of moments, is that one can use a pseudoinverse $M_{p}^{\dagger}$, $M_{p}M_{p}^{\dagger}=I$, according to
(3) $\overline{w}\leftarrow M_{p}^{\dagger}Mu,$
and moments are conserved as in the particle deposition [17, 34]. Thus, if $L$ conserves moments [16] this entire process of applying a continuum operator in a PIC method conserves moments.
2.2 Pseudoinverses and idempotent projections
There are two basic approaches to the pseudoinverse: an appropriate Krylov methods such as LSQR or Moore-Penrose. Both of these solvers are $l_{2}$ projections, but there are alternative norms such as $L_{2}$ that could be investigated. Moore-Penrose is attractive because it is easier to precondition a square matrix, especially for batch solvers [3]. Krylov methods are attractive as they can solve singular systems transparently. Preconditioning LSQR requires some effort, but the pseudoinverse solves in this work are very well conditioned and unpreconditioned LSQR works well. Both of these solvers are available in PETSc without explicitly constructing the normal equations using a shell matrix with the operator $M_{p}M_{p}^{T}$ in Moore-Penrose. Though we use LSQR in this work, the use of a Moore-Penrose pseudoinverse makes some of the analysis of stability clearer in §3.1 because it is defined with standard linear algebra operations: $M_{p}^{\dagger}\equiv M_{p}^{T}\left(M_{P}M_{p}^{T}\right)^{-1}$ and stability is easier to understand if we restrict ourselves to non-singular matrix solves.
If a collision operator is not used, $L=I$, then combining (2) and (3) results in the remapping algorithm
(4) $\overline{w}\leftarrow M_{p}^{T}\left(M_{P}M_{p}^{T}\right)^{-1}M_{p}w,$
which is a type of “coarse-graining” algorithm, a mechanism to add numerical entropy dissipation [6, 38].
Idempotent property of projections
Information is lost while projecting a particle representation of a function onto a FE basis if the number of particles exceeds the number of FE basis functions, which is typically the case of interest. An attractive property of (4) is that information is only lost on the first application in that
$\overline{\overline{w}}=M_{p}^{T}\left(M_{P}M_{p}^{T}\right)^{-1}M_{p}{\overline{w}}=M_{p}^{T}\left(M_{P}M_{p}^{T}\right)^{-1}M_{p}M_{p}^{T}\left(M_{P}M_{p}^{T}\right)^{-1}M_{p}w=M_{p}^{T}\left(M_{P}M_{p}^{T}\right)^{-1}M_{p}w={\overline{w}}.$
Thus $\overline{\overline{w}}={\overline{w}}$ and the coarse-graining operator is idempotent, which is an elegant property in that this process does not, in a sense, evolve the operator although it does add diffusion.
3 A particle resampling method
The key observation from this discussion is that in computing (4), the distribution function is entirely represented on the FE space after $M_{p}$ is applied to $w$ – particle weights and positions are no longer needed. A new set of particle positions, essentially any new set, can be created. A new particle mass matrix, $\overline{M}_{p}$, can be computed and (4) can be continued with $\overline{M}_{p}$. Moments are conserved because it is provable, and experimentally demonstrated, that the projection to the grid preserves moments and the projection from the grid preserves moments. The framework for this resampling method is to rearrange (4), by projecting back to a new set of particles after the deposition according to:
  • use (2) to deposit the distribution function on to the FE grid $c\leftarrow M_{p}w$,
  • create a new set of particles to generate a new particle mass matrix $\overline{M}_{p}$,
  • apply a pseudoinverse to compute weights for the new particles $\overline{w}\leftarrow\overline{M}_{p}^{\dagger}c$.
Field preservation with resampling: ${\overline{\rho}}=\rho$
An attractive property of this resampling is that the right hand side of the field solves, Poisson and Ampere’s law solves, or collision operators are not affected by the resampling:
$\overline{\rho}=M^{-1}{\overline{M}_{p}}\overline{w}=M^{-1}{\overline{M}_{p}}{\overline{M}_{p}}^{T}\left({\overline{M}_{p}}{\overline{M}_{p}}^{T}\right)^{-1}M_{p}w=M^{-1}M_{p}w=\rho.$
3.1 Moore-Penrose stability
Care must be taken in the explicit inverse of $M_{p}M_{p}^{T}$ as it can be singular from, for example, empty rows of $M_{p}$ if there are no particles in an FE basis function. Defining the particle grid is under the control of the algorithm, unlike in coarse-graining and field solves, and sufficient constraints must be understood. A necessary condition for stability of the pseudoinverses is that there does not exist a set of vertices whose union of support of associated FE basis functions contains less particles than the number of vertices in the set. This is not rigorous but comes from the intuition that no set of equations (rows of $M_{p}$) should have less non-empty columns than rows, otherwise the matrix locally singular. Further we find that more particles than basis functions are required for stability in, for example, the $1D$ periodic direction if, with degree $Q$ element (e.g., $Q=2$, a Q2 element), there are $Q$ particles per cell, which results in an equal number of particles and basis functions (equations or vertices). $M_{p}$ is square in this case, but we observe that $\left(M_{P}M_{p}^{T}\right)$ is singular except for the special case where the particles and mesh points are aligned as in the direct remap method in §5. This criterion is not practical to check with an arbitrary set of particles, but we believe that a simple criterion is robust in our experience although not tight. With tensor elements, we use $\left(Q+1\right)^{D}$ particles per cell. We are able to use $Q^{D}$ particles per cell in the doubly periodic case in the direct remap test in §5 because $\overline{M}_{p}=I$. One could imagine a tighter criteria where boundary cells have $Q^{D}+1$ particles and interior cells have $Q^{D}$, but we have not tested this.
4 Numerical methods and Landau damping test
Consider the classic one-dimensional plasma test, Landau damping, given by the initial state,
(5) $f(x,v)=\frac{1}{\sqrt{2\pi}}e^{-v^{2}}\left(1+A\cos\left(kx\right)\right),$
(6) $(x,v)=\left[0,2\pi/k\right]\times\left[-v_{max},v_{max}\right],$
where $k=0.5$, $v_{max}=6$, and we consider three values for the wave amplitude, $A$: $0.5$, $0.01$ and $0.0001$. The Landau damping test is a popular choice for Vlasov benchmarking because it involves a number of purely kinetic effects, such as phase mixing, and it has simple analytical solutions. A detailed description of the analytical solutions to the Landau damping problem can be found in [11]. To focus on the electrostatic kinetic effects, we ignore collisional dynamics and reduce the full VML equations to the collisionless, magnetic-free Vlasov-Poisson (VP) system.
The two cases where the wave amplitude, $A$, is $0.01$ and $0.5$ are referred to as the linear and nonlinear landau damping test cases, respectively [19, 7]. A third test case, $A=0.0001$, sits well within the linear regime and also common. In the linear case, the small field perturbation is damped out at a rate of $\gamma=-0.153$ in favor of a more homogeneous field. However, in collisionless tests the growth of subgrid modes can disrupt the field damping and cause large gradients to develop in the phase space. These large gradients lead to a sudden regrowth of the field. Previous work [19] has show that the inclusion of collisions can remove these subgrid modes, damping the field to machine precision with a continuum code.
In the nonlinear case, $A=0.5$, the field decay reverses much earlier and the general dynamics differ from that of the linear case. The two primary explanations for this earlier field resurgence are the stronger interaction between the potential well created by the electric field which resonates with and accelerates more of the particles, and the increased phase mixing, evident in phase-space diagrams. To fully capture these dynamics, the nonlinearized form of the Vlasov equation must be considered. Significant work was done by Villani and Mouhot in [29] to analyze the nonlinear Vlasov equation and show that, while the nonlinear dynamics present in this system lead to a weaker initial decay of the electric field, over long enough time scales, the field will damp out, as it does in the linear case. Thus, it is vital to develop the tools necessary to capture the long-time evolution of these kinetic plasmas structures. From [19] and [7], we expect the initial damping rate of the field to be $\gamma_{1}=-0.286$ which quickly turns into a field growth at a rate of $\gamma_{2}=0.086$. We will use these values to verify our tests in later sections.
4.1 PETSc test harness
The testing codes for this paper are build on the PETSc-PIC framework [11, 33, 34], a recently developed PIC toolkit in the PETSc [5, 9]. The PETSc-PIC framework primarily relies on two modules to drive forward the particle and finite element space. These modules are DMSwarm [28] and DMPLEX [22], respectively. DMSwarm provides a fully parallel solution for particle methods (e.g. DEM, SPH, EFG) and for particle-mesh methods (e.g. PIC, FLIP, MPM, GIMP) while DMPLEX provides generic unstructured mesh creation, manipulation and I/O.
The a finite element method (FEM) is used to solve the field equations at each timestep. The PETSc FEM framework abstracts the construction of the finite element using the Ciarlet triple [8], consisting of a mesh object (DMPLEX), a finite-dimensional function space (PetscSpace), and a dual space (PetscDualSpace). This is all handled by the PetscFE object and can be customized from the command line. In previous work [11], simple $H^{1}$ finite element spaces have been sufficient in capturing the short timescale linear plasma kinetics. Thus, we will continue the use of these $H^{1}$ spaces in this work.
Particle pushing for the VP system relies on the characteristics of linear hyperbolic Vlasov equation which may be derived by first written a simplified form of the Vlasov equation,
(7) $\frac{\partial f_{\alpha}}{\partial t}+\frac{\partial\vec{x}}{\partial t}\cdot\nabla_{x}f_{\alpha}+\frac{\partial\vec{v}}{\partial t}\cdot\nabla_{v}f_{\alpha}=0$
$\frac{\partial f_{\alpha}}{\partial t}+\mathbf{z}\cdot\nabla_{\mathbf{q}}f_{\alpha}=0,$
where $\mathbf{q}=(\mathbf{x},\mathbf{v})$ is the phase space variable and $\mathbf{z}=(\mathbf{v},-q_{e}\mathbf{E}/m)$ is the combined force. The force term $-q_{e}\mathbf{E}/m$ is independent of velocity, and therefore (7) may be written in the conservative form,
(8) $\frac{\partial f_{\alpha}}{\partial t}+\nabla_{\mathbf{q}}\cdot(\mathbf{z}f_{\alpha})=0.$
Given this new advective form of the Vlasov equation, we can rewrite the equation for the characteristics $\mathbf{Q}=(\mathbf{X},\mathbf{V})$,
(9) $\frac{d\mathbf{Q}}{dt}=\mathbf{z},$
which re-expressed with the original phase-space variables gives,
(10) $\frac{d\mathbf{X}}{dt}$ $=\mathbf{V},$
$\frac{d\mathbf{V}}{dt}$ $=-\frac{q_{e}}{m}\mathbf{E}.$
Since particles follow characteristics, the Vlasov equation in the particle basis becomes
(11) $\frac{d\mathbf{x}_{p}}{dt}$ $=\mathbf{v}_{p},$
$\frac{d\mathbf{v}_{p}}{dt}$ $=-\frac{q_{e}}{m}\mathbf{E}.$
Solving the characteristic equations is conducted in PETSc with the TS module using explicit symplectic integrators, a subclass of geometric integrators introduced by Ruth in [35]. In general, for PIC models, explicit integrators are not energy conservative and have a tendency to increase total energy over long time scales through “numerical heating”. In previous work [20, 11, 33], however, explicit symplectic integrators have been shown to achieve exact conservation of mass and momentum, as well as a stable approximate conservation of the system energy. The PETSc TS module contains well-tested implementations for first- to fourth-order symplectic integrators. For full Tokamak models and collisional cases, the TS module also contains a variety of implicit time integrators. These include the recently added discrete gradients method which has been tested on both VP and collisional Landau systems [12]. With these implicit methods, larger time steps can be taken while remaining stable, capturing long-time physical phenomena. Furthermore, exact energy conservation has been previously shown using implicit methods [27]. In this work, however, we are interested in capturing the fastest waves in the Landau damping system. Thus, explicit methods are more appropriate and less costly than implicit integrators. We choose a first-order symplectic integrator, called symplectic Euler.
4.2 Particle grids and finite element order
The phase space continuum grids in this work are on regular 90 degree lattice with an option for simple r-refinement in velocity space. The particle grids are defined with a Cartesian particle grid, of a fixed size, in each phase space cell, similar to Lapenta (Figure 1, [25]). For simplicity, the original grid is used for resampling in this work and adaptivity strategies are left for future work. As discussed in §3.1, with periodic boundary conditions in the spatial dimension and natural boundary conditions in velocity dimension we use at least $Q+1$ particles in each dimension in each phase space cell for stability of the pseudoinverse, but we found that using more particles per cell is often desirable for noise reduction.
The test harness is equipped with a simple r-adaptivity capability where points are pushed toward the origin in velocity space to better represent a Maxwellian distribution. Figure 1 shows the electric field (E) on uniform and r-refined versions of a $64\times 128$ particle grid $X\times V$, with particle clustering around $v=0$ and the initial perturbation in $x$ of the electric field.
y-
Figure 1: E field on $64\times 128$ particle grid (y-axis is velocity): uniform distribution (left); r-refinement (right)
5 Finite element version of the direct remap method
An approach similar to our method that aligns low-order continuum space points with the particle points allows for a direct map from the grid to particles. A cell-centered finite difference version of this approach was developed by Colella et al. on Cartesian grids [41, 30], and with phase-space adaptive mesh refinement (AMR) [31]. These experiments are intended to observe the qualitative effect of high-order grid to particle mapping. Quantitatively, the direct map approach conserve energy to only four digits while the energy is conserved to all double precision digits with $Q2$ in velocity (see data the src/A.01 directory for $Q1$ and in src/A.0001 for $Q2$ in the data repository). The pseudoinverse projection also allows, and in fact requires, more that one particle per phase-space cell, which is common in practice and allows for adjusting the smoothing of the distribution that reduces weight variance in a spatial cell, which is a common criterial for resampling algorithms [10].
Our test harness supports a direct map method in a vertex-centered, finite element context. Linear $Q1$ elements are used in the velocity and space dimension on Cartesian grids. Particles are placed at vertices and the domain is doubly periodic, which is immaterial given that there should be negligible density at the velocity boundary. The salient feature of this construct is that the particle mass matrix is the identity, $\overline{M}_{p}=I$, and the pseudoinverse vanishes.
Direct map resampling is tested with a linear Landau damping, $A=0.01$, problem with a $128\times 256$ particle grid and $V_{max}=6.0$ and results in Figure 2 agree well with amplitude reported in [30] (Figure 3.1).
Figure 2: Particle electric field amplitude with linear Landau damping, $A=0.01$, with no resampling, a direct remapping finite element version of Myers et al., and the projection method
$Q_{1}-Q_{1}$ is used for the projection method with one particle per cell, or vertex, to mimic the direct map method, and the projection method uses a $2\times 2$ particle grid per cell and half as many cells in each dimension to maintain the same number of particles for all three tests and to ensure stability of the pseudo-inverse. This test show a modest improvement in the behavior of the solver, which is simply an observation of the effect of the high-order mapping to back to particles (linear instead of the constant of the direct map). This data also shows the efficacy of resampling in that it suppresses the noise observed without resampling for the $A=0.01$ case.
6 Numerical experiments with pseudoinverse resampling
This section investigates results of the new projection resampling method on a linear and a nonlinear Landau damping tests. An additional algorithmic feature used in these experiments is to use an ad-hoc method to enforce continuous, $C^{0}$ electric field. One can use an $H(div)$ Poisson solver and this is a subject of future work, but in this work the electric field computed at particles, with a $C^{0}$ potential field, is not $C^{0}$, which can cause particles to artificially reflect off of cell boundaries and with the highly regular grids used here large instabilities result with $Q2$ elements. These instabolites are at least partialy due to lack of a quite start method in this work [36].
To fix this problem the electric field is projected to the vertices and then back to the particles. This method precludes energy conservation, but there are other sources of loss of energy conservation in our Vlasov-Poisson solver and the method is convergent. We have verify that the energy remains unchanged to machine precision from the full resampling algorithm by printing out the second moment before and after resampling and use a very tight tolerance in the pseudoinverse solve.
6.1 Nonlinear Landau damping, $A=0.5$
The nonlinear Landau damping test case, $A=0.5$, has been studied extensively in the literature (Kraus Table 5.1 tabulates several results of previous work [19]). Cheng and Knorr test with a $32\times 128$ cell continuum grid solver and a time step of $\frac{1}{8}$ that is not only quantitatively similar to our results (Figure 3(d)) but qualitatively similar (Figure 4 in [7] and Figure 5 in [21] shows $\gamma_{1}$ measured with the first to peaks, and the fourth peak while the third is low). Figure 3 show convergence studies on the amplitude of the electric field for time step, resampling rate and particle and continuum grid resolution.
(a) Particle number convergence
(b) Convergence in $\Delta t$
(c) Convergence in resampling period
(d) Highly converged
Figure 3: Converge study of nonlinear Landau damping, $A=0.5$
These convergence tests establish parameters for an highly resolved analysis in Figure 3(d) that agrees well with Cheng and Knorr: $\gamma_{1}=-0.281$ as compared to our rate of $-0.278$, and a growth rate of $0.084$ vs our rate of $0.091$. Figure 3(b) show convergence with time step. The differences are only visible in the rebound stage and approximately second order accuracy is observed. The modest effect of resampling in this nonlinear case is observed in Figure 3(c), where resampling appears to increase the amplitude in the rebound stage. In comparison with the (clean) continuum results of Kraus (Figure 5.11 [19]) the onset of rebound is about the same but slope of the rebound stage with our code, without resampling appears to be higher and this slope is $\gamma_{2}=0.07458$ in Kraus and $\gamma_{2}=0.07885$ in Figure 3(c).
6.2 Linear Landau damping
A linear case of Landau damping, $A=0.0001$, demonstrates the potential of the projection algorithm. Figure 4 show the electric field amplitude with a variety of resampling rates with $Q2$ spaces in both space and velocity.
Figure 4: $E_{max},A=0.0001$, long time simulation with and without resampling, $Q2-Q2$ spaces
No resampling is clearly very noisy and all of the resampling tests are free of noise, demonstrating the potential of this projection method for quiet, long time simulations. Resampling does seem to kick the plasma into growth phase faster, given that tests with a high resampling rate start to grow earlier. The damping rate of $\gamma_{1}=-0.15348$ agrees with theory, $\gamma_{1}=-0.153$, reported in [19].
7 Conclusion
The paper develops a new approach to particle resampling that uses a conservative projection, a pseudoinverse, to map any distribution of particles to essentially any other particle grid while conserving all moments up to the degree of polynomial that the projection function space can represent exactly. This method is evaluated with a simple projection grid, the original grid, on standard Landau damping problems. Tests where noise is problematic, the linear cases, show that resampling reduces noise considerably and coherent dynamics are maintained for long times, where as the solution becomes essentially all noise without resampling.
There area several areas of future work in developing adaptivity strategies for this method, for both particles and the projection grid, such as limiting the numerical entropy inherently generated in resampling, using adaptive continuum grid adaptivity (AMR) and using a modified original grid with particle splitting and coalescing techniques.
Entropy
With entropy measures, from our particle Landau collision operator [33], we can determine the continuum grids required for resampling or a continuum collision operator [3] to keep entropy generation by the projection well below the entropy generated in the collision operator.
AMR
AMR continuum grids in velocity space are available in PETSc [3], and regular particle grids on each cell of an adapted grid, like a cubed sphere, is a path for generating adapted particle grids through continuum mesh adaptation as is done in Myers et al. [31].
Splitting and coalescing
These ideas, developed in many particle resampling methods, would allow for an incremental modification of the particle mesh to minimize cost and perhaps impact the physics less. We do observe that resampling increases the onset of rebound in Figure 4, which may imply that the physics is disturbed by the resampling and should therefor be minimized.
Increasing relevance
Understanding the effects of resampling on physics, beyond conserving moments and other structure like entropy stability, requires experimentation, verification and validation, with more complex models like the Ion Temperature Gradient (ITG) instability.
Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences and Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) Partnership programs under Contract No. DE-AC02-05CH11231 at Lawrence Berkeley National Laboratory, the Office of Naval Research and by an appointment to the NRC Research Associateship Program at the U.S. Naval Research Laboratory, administered by the Fellowships Office of the National Academies of Sciences, Engineering, and Medicine. DSF and MGK were partially supported by NSF CSSI grant 1931524.
Appendix A Artifact description and reproducibility
PETSc output files with all data, provenance information, and reproducibility instructions for all tables and plots can be obtained from git@gitlab.com:markadams4/resampling-paper.git. This includes the python scrips that generates the plots and run scripts, makefiles and PETSc resource files used to generate the data, and the test harness code in src. The src/A.X directories has data for $A=0.$X. The exact PETSc versions (SHA1) are in the data files, with the provenance data, all parameters used in each test, but any PETSc version from v3.22 should suffice to reproduce this data.